OpenStudy (anonymous):

What is the sum of an 11–term arithmetic sequence where the first term is 5 and the last term is 55? 300 330 360 390

OpenStudy (gw2011):

The amswer is 330. Each number in the sequence increases by 5.

OpenStudy (anonymous):

Gauss' forrmula sum the first and last term of the series and divide by the number of pairs n/2

OpenStudy (anonymous):

\[s _{n}=\frac{n}{2}\left( a _{1}-a _{n} \right)\]\[s _{11}=\frac{11}{2}(5+55)=330\]

OpenStudy (anonymous):

@kiss.. think of it this way: suppose they were asking you to sum 11,000 terms, not 11 terms! the Gauss formula is easy to remember and useful